OpenStudy (anonymous):

A direct variation includes the points (10, 20) and (3, n) . Find n . Write and solve a direct variation equation to find the answer.

OpenStudy (anonymous):

use 10/20=3/n or 3/n=1/2 hence n=6

OpenStudy (anonymous):

Equation: 10/20 = 3/n, given you don't know n yet. Do you know what N equals now????